# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v1

@@ 5.3.1 Potential Benefits and Risks
For some individuals, particularly those with strong existing social networks and positive psychological profiles, companion chatbots may serve as a beneficial supplement to human social interaction. These users may experience higher social confidence in parallel with their chatbot use. Our data suggest that neurotic individuals, who tend to be lonely and socially isolated, may experience reduced loneliness when spending more time with chatbots, but previous research suggests that more socially isolated individuals are at greater risk of problematic use of online services [Reer et al.(2019), O’Day and Heimberg(2021)]. As our model should be interpreted cautiously, the risk of chatbots replacing human interactions and potentially exacerbating loneliness and social withdrawal should be considered.
5.3.2 Personalized Approaches
The diverse user profiles identified in our study suggest a need for personalized approaches in chatbot design and deployment. Chatbots could potentially be tailored to different user types, with features and interaction styles that complement rather than replace human social interaction for vulnerable users.
5.3.3 Chatbots as Loneliness Interventions
For individuals with limited access to human social support, companion chatbots could potentially serve as an intervention to reduce loneliness and its associated health risks. However, careful consideration must be given to how such interventions are designed and implemented to ensure they support, rather than hinder, the development of human social connections. Additionally, the direction of causality between loneliness and companion chatbot usage is not clear in our data; further work should be done to investigate the relationship further.
5.3.4 Ethical Considerations & Long-term Societal Impacts
@@
The potential for some users to engage in deep, emotional relationships with chatbots raises ethical concerns. Developers and policymakers need to consider how to balance the potential benefits of AI companionship with the risks of over-reliance on artificial relationships. This may involve implementing safeguards or guidance systems within chatbots to encourage healthy usage patterns and human social interaction.
@@
As AI technology continues to advance and companion chatbots become more sophisticated, it is crucial to consider the long-term societal impacts of widespread AI companionship. Questions arise about how these technologies might shape social norms, expectations for human relationships, and the development of social skills, particularly for younger generations. Should we as a society push for less artificial companionship, or should we encourage a healthy mix of artificial and real companionship? Consider the case of a profoundly lonely and emotionally vulnerable individual. They might cling to a relationship that can provide any crumb of affection, no matter how unhealthy it becomes. Could companion chatbots fulfill that desperate need in a beneficial way, even if “unhealthy attachment” to the chatbot is part of the process? Perhaps alleviating loneliness enough to reduce the individual’s vulnerability could allow them to become more conscientious about seeking out and maintaining healthy, respectful human relationships.
@@ 5.4 Future Research Directions
Our study opens up several avenues for future research that could further elucidate the relationship between companion chatbot usage and loneliness. Longitudinal studies would be particularly valuable in understanding the causal relationships between chatbot usage, loneliness, and social outcomes over time. Such studies could track changes in social behaviors, relationship patterns, and psychological well-being as individuals engage with companion chatbots over extended periods.
@@
Focused studies on specific user groups, such as adolescents, older adults, or individuals with social anxiety, could yield valuable insights for developing targeted designs or guidelines.
As companion chatbots become more prevalent, it will be crucial to investigate their long-term effects on social skills, relationship expectations, and overall social functioning. This could involve extended longitudinal studies or even intergenerational research to understand how exposure to AI companionship from a young age might shape social development and interpersonal dynamics.
@@ 5.4.1 Limitations and Considerations
We note a number of limitations in our study.
Survey design and recruitment. While our adaptations of established scales were based on factor loadings from previous studies and tailored to our research context, we acknowledge that these modifications may affect the scales’ validated psychometric properties. Future work should consider validating these adapted scales specifically for chatbot-related research. Due to the extensive nature of our survey, we prioritized reducing participant fatigue over maintaining the full length of each scale. This trade-off allowed us to explore a broader range of factors but may have impacted the depth of measurement for individual constructs. Additionally, our participants were primarily from the United States; applying our findings to other demographics should be done cautiously.
Clustering. We note that our clusters have their own set of limitations. While the number of clusters has statistical grounding, the final number can be considered somewhat arbitrary and based on our subjective judgments of interpretability. The heterogeneity of covariance matrices indicated by Box’s M-test requires caution in interpreting these results. Even so, this 6-cluster solution provides a starting point for further work on styles of use.
Regression Model. The assumption of linear relationships may not capture complex nonlinear relationships in the population of users of companion chatbots. Furthermore, this model is based on correlation and does not determine causality. Future research should explore potential non-linear relationships and employ longitudinal designs to better understand the causal mechanisms at play. The exploratory nature of our process makes much room for error, requiring more in-depth research to understand the relationship between companion chatbot usage and loneliness.
@@ 6 Conclusion
This study examines the relationship between companion chatbot usage and loneliness, offering insights into the complex interplay of psychological, social, and technological factors that shape this relationship. Our findings highlight the diverse ways in which individuals engage with and are affected by companion chatbots, ranging from socially fulfilled dependent users who experience enhanced social confidence to lonely dependent users at risk of social withdrawal.
The model we developed through regression analysis reveals that while chatbot usage does not directly predict loneliness, it interacts with factors such as social attraction to close others and neuroticism in complex ways. This suggests that the relationship between companion chatbot usage and loneliness is not uniform but highly dependent on individual characteristics and social contexts. Our cluster analysis further underscores this complexity, identifying seven distinct user profiles that differ not only in their patterns of chatbot usage but also in their psychological characteristics and social outcomes. These profiles suggest that different types of users may benefit from or be at risk due to companion chatbot use. Our work underscores the need for personalized approaches in chatbot design and deployment.
As AI companions become increasingly prevalent, it is crucial to consider their potential long-term impact on social norms, relationship expectations, and the development of social skills. The increasing usage of companion chatbots in society raises ethical questions about the role of AI in addressing loneliness and emotional vulnerability.
As we continue to navigate the evolving landscape of human-AI interaction, further research is needed to fully understand the causal relationships between chatbot usage and social outcomes, explore potential non-linear effects, and investigate long-term impact across different cultural contexts and user groups. By pursuing these research directions, we can work towards harnessing the potential benefits of AI companionship while mitigating risks, ultimately supporting the development of technologies that genuinely enhance human well-being and social connection in our increasingly digital world.
7 Acknowledgements
@@ Appendix A Significant Correlations with Key Variables
Below, we include a list of significant (corrected $p<0.05$) correlations with chatbot usage frequency score, chatbot session length score, and loneliness (ULS-8). Negative correlations have text colored in blue.
@@ Appendix B Cluster Profiles
A heatmap of cluster means for variables at their original scales in each cluster, colored based on their deviation from the mean, can be seen in Figure 17.

### v2

@@ 5.3.1 Theoretical Contributions
Potential Benefits and Risks. For some individuals, particularly those with strong existing social networks and positive psychological profiles, companion chatbots may serve as a beneficial supplement to human social interaction. These users may experience higher social confidence in parallel with their chatbot use. Our data suggest that neurotic individuals, who tend to be lonely and socially isolated, may experience reduced loneliness when spending more time with chatbots, but previous research suggests that more socially isolated individuals are at greater risk of problematic use of online services [Reer et al.(2019), O’Day and Heimberg(2021)]. As our model should be interpreted cautiously, the risk of chatbots replacing human interactions and potentially exacerbating loneliness and social withdrawal should be considered.
Implications for HCI Design and Practice. Our findings on problematic use suggest the need for built-in mechanisms to detect patterns of potentially harmful usage, provide user feedback about usage patterns, and implement interventions when problematic patterns emerge. Additionally, offering resources for maintaining healthy human relationships alongside chatbot use may prove beneficial if they can help improve the user’s social support.
Further, if the user’s traits can be collected, designing adaptive interfaces for social chatbots based on user traits to encourage or discourage certain uses may be an avenue for mitigating risk of negative effects for at-risk individuals (such as those with high neuroticism and signs of problematic use). Chatbots could potentially be tailored to different user types, with features and interaction styles that complement rather than replace human social interaction for vulnerable users.
For individuals with limited access to human social support, companion chatbots could serve as an intervention to reduce loneliness and its associated health risks. However, the design and implementation of these interventions must be carefully considered. Future HCI research should consider multiple pathways of influence when studying and designing AI companion technologies.
@@
Ethical Considerations & Long-term Societal Impacts. The potential for some users to engage in deep, emotional relationships with chatbots raises ethical concerns. Developers and policymakers need to consider how to balance the potential benefits of AI companionship with the risks of over-reliance on artificial relationships. This may involve implementing safeguards or guidance systems within chatbots to encourage healthy usage patterns and human social interaction.
@@
As AI technology advances and companion chatbots become more sophisticated, it is crucial to consider the long-term societal impacts of widespread AI companionship. Questions arise about how these technologies might shape social norms, expectations for human relationships, and the development of social skills, particularly for younger generations. Should we as a society push for less artificial companionship, or should we encourage a healthy mix of artificial and real companionship? Consider the case of a profoundly lonely and emotionally vulnerable individual. They might cling to a relationship that can provide any crumb of affection, no matter how unhealthy it becomes. Could companion chatbots fulfill that desperate need in a beneficial way, even if “unhealthy attachment” to the chatbot is part of the process? Perhaps alleviating loneliness enough to reduce the individual’s vulnerability could allow them to become more conscientious about seeking out and maintaining healthy, respectful human relationships.
@@ 5.4 Future Research Directions
Several key areas warrant further investigation. Longitudinal studies would be particularly valuable in understanding the causal relationships between chatbot usage, loneliness, and social outcomes over time. Such studies could track changes in social behaviors, relationship patterns, and psychological well-being as individuals engage with companion chatbots over extended periods.
@@
Focused studies on specific user groups, such as the clusters we identified or specific demographic groups such as adolescents, older adults, or individuals with social anxiety, could yield valuable insights for developing targeted designs or guidelines.
As companion chatbots become more prevalent, it will be crucial to investigate their long-term effects on social skills, relationship expectations, and overall social functioning.
@@ 5.4.1 Limitations and Considerations
Survey design and recruitment. As our survey consists entirely of self-reported measures for usage characteristics and psychosocial measures, we face a considerable risk of recall bias. Our custom instruments for usage characteristics offered fairly broad options to reduce potential inaccuracies, though this also makes them coarse measures. Conducting a study involving monitoring or reporting of actual use would provide more specific and accurate usage data, and is in our plans for future work.
While our adaptations of established scales were based on factor loadings from previous studies and tailored to our research context, we acknowledge that these modifications may affect the scales’ validated psychometric properties. Future work should consider validating these adapted scales specifically for chatbot-related research. Due to the extensive nature of our survey, we prioritized reducing participant fatigue over maintaining the full length of each scale. This trade-off allowed us to explore a broader range of factors but may have impacted the depth of measurement for individual constructs. Additionally, our participants were primarily white/Caucasian individuals from the United States; applying our findings to other demographics should be done cautiously.
Clustering. We note that our clusters have their own set of limitations. While the number of clusters has statistical grounding, the final number can be considered somewhat arbitrary and based on our subjective judgments of interpretability. The heterogeneity of covariance matrices indicated by Box’s M-test requires caution in interpreting these results. Even so, this 7-cluster solution provides a starting point for further work on styles of use.
Regression Model. The exploratory nature of our model selection process warrants cautious interpretations and further investigation into the relationships it implies. The assumption of linear relationships may not capture complex nonlinear relationships in the population of users of companion chatbots. Further, due to the numerous predictors involved in the exploration and the final model, the risk of Type II errors is considerable [Lavery et al.(2019)]; our analysis may not capture all potential relationships of these variables.
Additionally, while we can theorize about the directions of relationships between variables, our cross-sectional design limits causal inferences. Future research employing experimental designs or longitudinal studies could provide more definitive evidence of causal relationships and explore potential non-linear effects. We plan to follow up with our participants to see changes in psychological, social, and usage characteristics at different time points.
@@ 6 Conclusion
This study examines the relationship between companion chatbot usage and loneliness, developing a model that explains approximately 50% of variance in loneliness through the complex interplay of psychological, social, and technological factors. Our findings reveal that while chatbot session length does not directly predict loneliness, this relationship is mediated by problematic use and moderated by social attraction to close others, highlighting the nuanced ways in which AI companions influence social well-being.
The model we developed through regression analysis identifies key mechanisms, including problematic use as a mediator and social attraction as a moderator, while accounting for personality traits such as neuroticism and social network characteristics. This suggests that the impact of companion chatbot usage on loneliness operates through complex pathways rather than direct effects.
Our cluster analysis further underscores this complexity, identifying seven distinct user profiles that demonstrate how similar usage patterns can lead to markedly different outcomes. For instance, socially fulfilled dependent users maintain healthy relationships despite intensive chatbot use, while lonely moderate users with similar usage patterns show signs of social withdrawal. These contrasting outcomes emphasize the importance of individual characteristics and social contexts in shaping the effects of AI companionship.
As AI companions become increasingly prevalent, these findings have crucial implications for design and deployment. The identification of distinct user profiles and key mediating factors suggests the need for personalized approaches and built-in mechanisms to detect and intervene in potentially harmful usage patterns. The contrasting outcomes between different user types raise important questions about how to design AI companions that support psychosocial well-being.
Future research should explore the causal relationships and long-term impacts of AI companionship through longitudinal studies while investigating how different design approaches might better support healthy engagement patterns. As we continue to navigate the evolving landscape of human-AI interaction in an increasingly digital world, understanding these complex relationships becomes crucial for developing technologies that genuinely enhance human well-being.
7 Acknowledgments
@@ Appendix A Significant Correlations with Key Variables
Below, we include a list of significant (corrected $p<0.05$) correlations with chatbot usage frequency score, chatbot session length score, and loneliness (ULS-8). Negative correlations have text colored in blue. Additionally, the correlation matrix with original variable names is seen in Figure 18.
Figure 18: Correlation matrix of psychological, social, and usage characteristics with original variable names.
@@ Appendix B Cluster Profiles
A heatmap of cluster means for variables at their original scales in each cluster, colored based on their deviation from the mean, can be seen in Figure 19. Figure 20 is a demographic table with percentages of each category for each cluster. Detailed descriptions for each cluster can be found below.
@@
Figure 20: Demographic table for study participants, including overall percentage for each category and each cluster.
B.0.1 Cluster 0: Disengaged Light Users
B.0.2 Cluster 1: Well-Adjusted Moderate Users
B.0.3 Cluster 2: AI-Wary Light Users
B.0.4 Cluster 3: Lonely Moderate Users
B.0.5 Cluster 4: Fulfilled Dependent Users
B.0.6 Cluster 5: Lonely Light Users
B.0.7 Cluster 6: Socially Challenged Frequent Users